# StoxLSTM: A Stochastic Extended Long Short-Term Memory Network for Time Series Forecasting

### v1

@@ 4.1.5 Results
Table 2: Long-term time series forecasting results. The best results are in bold and the second best are in underline.
@@
In long-term forecasting tasks, xLSTM-based models generally demonstrate superior performance compared to Transformer-based and Linear-based models, highlighting the strength of recurrent architectures like xLSTM in capturing long-range temporal dependencies. Among these, StoxLSTM achieves the best results. For instance, on the Weather dataset, StoxLSTM attains an average MSE of 0.161 across different prediction horizons, which is 26.5% lower than the xLSTM-Mixer’s average MSE of 0.219. Similarly, on the Electricity dataset, StoxLSTM achieves an average MSE of 0.139, outperforming xLSTM-Mixer’s 0.153 by 9.1%, further demonstrating its consistent advantage. Notably, StoxLSTM maintains this leading performance across other datasets such as Solar, ETTh1, and ETTm1, reflecting its robustness and generalizability in diverse long-term forecasting scenarios.
@@
In short-term forecasting, although Linear-based models show competitive results, StoxLSTM still surpasses them by a significant margin. Specifically, across four traffic datasets (PEMS03, PEMS04, PEMS07, and PEMS08), StoxLSTM achieves an average MAE of 12.81, outperforming the best Linear-based model, TimeMixer++, which records an MAE of 15.91, representing a 19.49% improvement. Furthermore, StoxLSTM attains an average MAPE of 8.24, which is 18.25% lower than TimeMixer++’s 10.08. In terms of RMSE, StoxLSTM achieves 19.98 on average, outperforming TimeMixer++’s 27.03 by 26.08%. These consistent improvements across multiple metrics and datasets underscore StoxLSTM’s effectiveness in capturing short-term temporal patterns with high accuracy. Some of the forecasting results are shown in Figs. 4 to 6.
Figure 4: The real and forecasting sequences of the last 6 out of 321 time series from the electricity dataset are shown. The orange line denotes the forecasted values generated by StoxLSTM, while the blue line represents the ground truth. The prediction horizon is set to 336 time steps.
@@
Figure 5: The real and forecasting sequences of the last 6 out of 862 time series from the Traffic dataset are shown. The orange line denotes the forecasted values generated by StoxLSTM, while the blue line represents the ground truth. The prediction horizon is set to 96 time steps.
@@
Figure 6: The real and forecasting sequences of the last 6 out of 170 time series from the electricity dataset are shown. The orange line denotes the forecasted values generated by StoxLSTM, while the blue line represents the ground truth. The prediction horizon is set to 48 time steps.
@@
Overall, these results demonstrate that StoxLSTM consistently delivers state-of-the-art accuracy across both short- and long-term forecasting tasks, validating its effectiveness and robustness in diverse time series forecasting scenarios. Supplementary experimental details are provided in B.2.
@@ 4.2 Model Analysis
Experimental results clearly demonstrate that all components—series decomposition, patching, channel independence, stochastic latent variables, and the SSM structure—contribute positively to the overall performance of StoxLSTM. Focusing first on the SSM modeling aspect, both the original StoxLSTM and its variants without stochastic latent variables (w/o Stochastic) and without series decomposition (w/o SD) consistently outperform P-sLSTM across all experiments, achieving lower MSE and MAE. This confirms that the state space model enhancement to the xLSTM architecture plays a crucial role in improving forecasting performance. Moreover, compared to the w/o Stochastic variant, StoxLSTM with stochastic latent variables generally achieves better and more stable prediction results, indicating that modeling latent states as random variables better captures the underlying temporal dependencies in time series data and enhances robustness. In contrast, the model variant without patching and channel independence exhibits notable instability: while it achieves excellent results in certain scenarios, its performance deteriorates significantly in others. This suggests that patching and channel independence not only improve robustness but also enable the model to effectively handle longer look-back lengths under constrained computational and memory resources, thereby reducing the risk of overfitting.
@@
Figure 7: Heatmaps of four tensor groups—$x$, $x_{p}$, $z$, and hh—from StoxLSTM when predicting a specific instance in the Electricity dataset. Different colors indicate different values, with brighter colors corresponding to larger values.
Within the inference model, values of $x_{p1:pN}$ and $g_{1:N}$ exhibit limited variation across patches but display noticeable differences across dimensions within each patch. The latent states effectively capture this pattern: both $z\_mean_{1:N}$ and $z\_logvar_{1:N}$ remain relatively stable across patches while preserving significant variability across dimensions. This suggests that the latent representation encodes consistent temporal patterns while retaining important feature-level distinctions.
Notably, through variational inference, the latent states learned by the generative model similarly reflect these characteristics—showing near-uniformity across patches alongside meaningful variation across dimensions within each patch. Such a well-structured latent representation, combined with the hidden states $h_{1:N}$, offers a comprehensive and informative summary of the time series, thereby supporting accurate and robust forecasting.
As demonstrated in Fig. 7, the variational inference training converges effectively, and the latent states successfully capture the underlying temporal features of the data. This visualization not only validates the model’s ability to learn meaningful representations but also highlights the interpretability of the latent space in modeling complex temporal dynamics.
@@ 5 Conclusion
In this paper, we propose StoxLSTM, a novel stochastic variant of the original xLSTM architecture. By integrating a designed SSM, StoxLSTM effectively captures the dynamic evolution of time series data. Our model demonstrates a strong capability in both long- and short-term forecasting tasks. Extensive experiments on multiple publicly available benchmark datasets show that StoxLSTM consistently outperforms existing state-of-the-art methods in terms of prediction accuracy and robustness. These results validate the effectiveness of combining the xLSTM framework with the SSM design to enhance temporal modeling. Future work will focus on further optimizing the model architecture and improving computational efficiency.
6 Acknowledgement
This work was supported by the National Natural Science Foundation (NSFC) of China under Grant 62301031.
@@ Appendix A ELBO Derivation of StoxLSTM
Appendix B Supplementary Experiments
B.1 CRPS Evaluation Metric
Generative time series forecasting can also use the Continuous Ranked Probability Score (CRPS) to evaluate the prediction results. CRPS is computed by comparing the ground truth with the cumulative distribution function (CDF):
[…]
For point forecasting models, the point prediction can be treated as a degenerate CDF. Substituting into the formula and simplifying yields:
[…]
indicating that CRPS for point forecasts is equivalent to MAE.
Table 5: Long-term time series forecasting results. The best results are in bold and the second best are in underline. The evaluation metric is CRPS.
Weather
96 0.167 0.184 0.187 0.208 0.214 0.198 0.296 0.384 0.336 0.237 0.205
192 0.206 0.226 0.236 0.256 0.254 0.241 0.336 0.544 0.367 0.282 0.245
336 0.225 0.266 0.272 0.297 0.296 0.282 0.380 0.523 0.395 0.319 0.265
720 0.254 0.323 0.326 0.350 0.347 0.334 0.428 0.741 0.428 0.362 0.334
Electricity
96 0.223 0.218 0.221 0.226 0.240 0.222 0.308 0.368 0.317 0.237 0.222
192 0.242 0.235 0.243 0.243 0.253 0.240 0.315 0.386 0.334 0.249 0.235
336 0.253 0.250 0.259 0.262 0.269 0.259 0.329 0.394 0.338 0.267 0.245
720 0.270 0.276 0.276 0.293 0.317 0.290 0.355 0.439 0.331 0.301 0.310
Solar
96 0.178 0.313 0.306 0.232 0.237 0.224 0.383 0.247 0.446 0.377 0.231
192 0.212 0.328 0.331 0.241 0.261 0.251 0.364 0.251 0.561 0.397 0.263
336 0.223 0.330 0.336 0.248 0.273 0.252 0.397 0.287 0.588 0.415 0.269
720 0.241 0.324 0.324 0.249 0.275 0.269 0.402 0.301 0.633 0.412 0.270
ETTh1
96 0.396 0.386 0.395 0.405 0.405 0.400 0.419 0.713 0.459 0.399 0.403
192 0.407 0.417 0.416 0.431 0.436 0.429 0.448 0.792 0.482 0.416 0.441
336 0.412 0.429 0.437 0.458 0.458 0.440 0.465 0.809 0.496 0.443 0.434
720 0.421 0.448 0.465 0.512 0.491 0.468 0.507 0.865 0.512 0.490 0.451
ETTh2
96 0.305 0.329 0.333 0.378 0.349 0.337 0.388 1.525 0.397 0.353 0.328
192 0.337 0.375 0.378 0.442 0.400 0.382 0.439 1.931 0.452 0.418 0.379
336 0.354 0.401 0.403 0.494 0.432 0.384 0.487 1.835 0.486 0.465 0.398
720 0.382 0.424 0.430 0.669 0.445 0.422 0.474 1.625 0.511 0.551 0.415
ETTm1
96 0.307 0.328 0.335 0.343 0.368 0.346 0.419 0.571 0.475 0.343 0.334
192 0.336 0.354 0.361 0.369 0.391 0.370 0.441 0.669 0.496 0.365 0.362
336 0.356 0.374 0.379 0.391 0.420 0.392 0.459 0.871 0.537 0.386 0.391
720 0.384 0.407 0.411 0.424 0.459 0.420 0.490 0.823 0.561 0.421 0.423
ETTm2
96 0.222 0.244 0.250 0.302 0.264 0.256 0.287 0.453 0.339 0.260 0.245
192 0.256 0.285 0.288 0.376 0.309 0.296 0.328 0.563 0.340 0.303 0.291
336 0.283 0.322 0.322 0.438 0.348 0.329 0.366 0.887 0.372 0.342 0.343
720 0.347 0.377 0.380 0.482 0.407 0.385 0.415 1.388 0.419 0.421 0.399
Traffic
96 0.257 0.236 0.242 0.332 0.268 0.249 0.366 0.391 0.388 0.282 0.253
192 0.310 0.241 0.253 0.342 0.276 0.256 0.373 0.379 0.382 0.287 0.258
336 0.320 0.250 0.261 0.345 0.283 0.264 0.383 0.420 0.337 0.296 0.263
720 0.335 0.283 0.287 0.357 0.302 0.286 0.382 0.472 0.408 0.315 0.282
ILI
24 0.579 0.707 0.694 1.309 0.883 0.754 1.260 1.449 1.182 1.081 0.823
36 0.680 0.712 0.722 1.330 0.870 0.870 1.080 1.463 1.038 0.963 0.835
48 0.710 0.721 0.725 1.326 0.815 0.815 1.078 1.542 1.145 1.024 0.818
60 0.768 0.737 0.715 1.384 0.788 0.788 1.157 1.543 1.114 1.096 0.839
B.2 Additional Experimental Results
We evaluate StoxLSTM’s CRPS across multiple datasets and prediction horizons in long-term forecasting tasks. For point prediction models, MAE and CRPS results coincide. We also assess CRPS for all baseline models. The results in Table 5 demonstrate that StoxLSTM achieves the lowest CRPS on most datasets, with xLSTM-Mixer ranking second. For instance, on the Solar dataset, StoxLSTM attains an average CRPS of 0.2135, which is 34.0% lower than the 0.3238 average CRPS of xLSTM-Mixer. Similarly, on the ETTm2 dataset, StoxLSTM achieves an average CRPS of 0.277, outperforming xLSTM-Mixer’s 0.307 by 9.8%. These results indicate that StoxLSTM provides more accurate probabilistic forecasts across diverse datasets.

### v2

@@ 4.1.3 Results
Table 2: Long-term time series forecasting results. The best results are in bold, and the second best are in underline.
@@
In long-term forecasting tasks, xLSTM-based models generally demonstrate superior performance compared to Transformer-based and Linear-based models, highlighting the strength of recurrent architectures like xLSTM in capturing long-range temporal dependencies. Among these, StoxLSTM achieves the best results. For instance, on the Weather dataset, StoxLSTM attains an average MSE of 0.151 across different prediction horizons, which is 31.1% lower than the xLSTM-Mixer’s average MSE of 0.219. Similarly, on the Electricity dataset, StoxLSTM achieves an average MSE of 0.139, outperforming xLSTM-Mixer’s 0.153 by 9.1%, further demonstrating its consistent advantage. Notably, StoxLSTM maintains this leading performance across other datasets such as Solar, ETTh1, and ETTm1, reflecting its robustness and generalizability in diverse long-term forecasting scenarios.
@@
In short-term forecasting, although Linear-based models show competitive results, StoxLSTM still surpasses them by a significant margin. Specifically, across four traffic datasets (PEMS03, PEMS04, PEMS07, and PEMS08), StoxLSTM achieves an average MAE of 12.81, outperforming the best Linear-based model, TimeMixer++, which records an MAE of 15.91, representing a 19.49% improvement. Furthermore, StoxLSTM attains an average MAPE of 8.24, which is 18.25% lower than TimeMixer++’s 10.08. In terms of RMSE, StoxLSTM achieves 19.98 on average, outperforming TimeMixer++’s 27.03 by 26.08%. These consistent improvements across multiple metrics and datasets underscore StoxLSTM’s effectiveness in capturing short-term temporal patterns with high accuracy.
Figs. 4 to 7 present partial visualizations of StoxLSTM’s prediction results. It can be observed that the prediction curves generated by StoxLSTM tend to be relatively smooth, and its ability to capture fine-grained variations in time series appears somewhat limited. This phenomenon could potentially be ascribed to the fact that StoxLSTM adopts a variational inference framework analogous to that of VAEs [18, 13] for learning the characteristics of latent variables.
(a) Horizon: 48
(b) Horizon: 96
(c) Horizon: 192
(d) Horizon: 336
@@
Figure 4: Visualization of the prediction results from StoxLSTM with different prediction horizons H={48, 96, 192, 336} from the Traffic dataset. The orange line denotes the prediction values generated by StoxLSTM, while the blue line represents the ground truth.
(a) Horizon: 96
(b) Horizon: 192
(c) Horizon: 336
(d) Horizon: 720
@@
Figure 5: Visualization of the prediction results from StoxLSTM with different prediction horizons H={96, 192, 336, 720} from the Electricity dataset. The orange line denotes the prediction values generated by StoxLSTM, while the blue line represents the ground truth.
(a) Horizon: 96
(b) Horizon: 192
(c) Horizon: 336
(d) Horizon: 720
Figure 6: Visualization of the prediction results from StoxLSTM with different prediction horizons H={96, 192, 336, 720} from the Solar dataset. The orange line denotes the prediction values generated by StoxLSTM, while the blue line represents the ground truth.
(a) Horizon: 24
(b) Horizon: 36
(c) Horizon: 48
(d) Horizon: 60
Figure 7: Visualization of the prediction results from StoxLSTM with different prediction horizons H={24, 36, 48, 60} from the ILI dataset. The orange line denotes the prediction values generated by StoxLSTM, while the blue line represents the ground truth.
@@
Overall, these results demonstrate that StoxLSTM consistently delivers state-of-the-art accuracy across both short- and long-term forecasting tasks, validating its effectiveness and robustness in diverse time series forecasting scenarios.
@@ 4.2 Model Analysis
Experimental results clearly demonstrate that all components—series decomposition, patching, channel independence, stochastic latent variables, and the SSM structure—contribute positively to the overall performance of StoxLSTM. Focusing first on the SSM modeling aspect, both the original StoxLSTM and its variants without stochastic latent variables (w/o Stochastic) and without series decomposition (w/o SD) consistently outperform P-sLSTM across all experiments, achieving lower MSE and MAE. This confirms that the state space model enhancement to the xLSTM architecture plays a crucial role in improving forecasting performance. Moreover, StoxLSTM with stochastic latent variables generally achieves better and more stable prediction results. Notably, the Solar-Energy and ETTm1 datasets exhibit higher randomness, with forecastability values [35, 14] of 0.33 and 0.46, respectively, while the Electricity dataset shows lower randomness, with a forecastability of 0.77. The performance improvements attributed to stochastic latent variables are more pronounced on the Solar-Energy and ETTm1 datasets compared to the Electricity dataset. This demonstrates that modeling latent states as stochastic variables effectively captures the inherent randomness in time series, enabling better forecasting accuracy and robustness, particularly for datasets with stronger stochasticity. This indicates that modeling latent states as stochastic variables effectively captures the inherent randomness in time series, leading to enhanced forecasting accuracy and robustness, especially for datasets with stronger stochasticity. In contrast, the model variant without patching and channel independence exhibits notable instability: while it achieves excellent results in certain scenarios, its performance deteriorates significantly in others. This suggests that patching and channel independence not only improve robustness but also enable the model to effectively handle longer look-back lengths under constrained computational and memory resources, thereby reducing the risk of overfitting.
4.2.2 Hyperparameter sensitivity
Patch size
We conducted a study on the impact of different patch sizes (16, 32, 48, 56, 64) on prediction performance across three datasets: Weather, ETTh2, and ETTm2. The results are presented in Figs. 8 to 10.
Figure 8: Impact of different patch sizes on prediction performance on the Weather dataset. Colors correspond to different prediction horizons, and the black line represents the average MSE or MAE.
Figure 9: Impact of different patch sizes on prediction performance on the ETTh2 dataset. Colors correspond to different prediction horizons, and the black line represents the average MSE or MAE.
Figure 10: Impact of different patch sizes on prediction performance on the ETTm2 dataset. Colors correspond to different prediction horizons, and the black line represents the average MSE or MAE.
Different datasets and prediction horizons may favor different patch sizes. As shown in Fig. 8, on the Weather dataset, patch sizes of 32 and 64 achieved the best prediction performance, yielding the lowest MAE and MSE values. Similarly, Fig. 9 demonstrates that a patch size of 32 leads to optimal results on the ETTh2 dataset. For the ETTm2 dataset, Fig. 10 indicates that the best performance is obtained when the patch size ranges between 48 and 56.
These findings suggest that the optimal patch size is dataset-dependent and may also vary with the prediction horizon. Nevertheless, patch sizes of 32 or 64 generally provide robust predictive accuracy across different scenarios. Notably, since the ETTm2 dataset has a higher sampling frequency compared to ETTh2, it likely requires a larger patch size to effectively capture the finer-grained temporal dynamics and achieve superior forecasting performance.
Initialization Method for $\bm{z_{0}}$
The initialization method for the initial latent state $\bm{z_{0}}$ in the SSM can also affect prediction performance. We compared two different initialization methods for $\bm{z_{0}}$—zero initialization (Zero) and Gaussian random initialization (Randn)—and evaluated their impact on prediction performance across different forecasting horizons on the ILI and ETTh2 datasets. The results are presented in Figs. 11 and 12.
Figure 11: Impact of different initialization methods for $\bm{z_{0}}$ on prediction performance under the ILI dataset.
Figure 12: Impact of different initialization methods for $\bm{z_{0}}$ on prediction performance under the ETTh2 dataset.
As illustrated in the figures, the choice of $\bm{z_{0}}$ initialization method leads to observable differences in prediction results. On the ILI dataset, Gaussian random initialization achieves an average MSE of 1.203 and an average MAE of 0.682 across different forecasting horizons, slightly outperforming zero initialization, which yields an average MSE of 1.201 and an average MAE of 0.688. Similarly, for the ETTh2 dataset, Gaussian random initialization attains an average MSE of 0.272 and an average MAE of 0.344, compared to zero initialization with an average MSE of 0.276 and an average MAE of 0.347.
Overall, Gaussian random initialization of $\bm{z_{0}}$ tends to yield better prediction performance. This may be because a randomly initialized latent state better captures the inherent stochasticity of the time series, thereby enhancing the model’s predictive accuracy and robustness.
@@ 4.2.2 Hyperparameter sensitivity
Figure 13: Heatmaps of five tensor groups—$x$, $\bm{x_{p}}$, $\bm{z}$, $\bm{h}$, and $\bm{g}$—from StoxLSTM when predicting a specific instance in the Electricity dataset. Here, $\bm{h}$ represents the output of the StoxLSTM recurrent unit during the generative process, while $\bm{g}$ denotes the output of the bidirectional StoxLSTM recurrent unit during the inference process. Different colors correspond to different values, with brighter colors indicating larger magnitudes.
Within the inference model, $\bm{x}_{p1:pN}$ and $\bm{g}_{1:N}$ vary little across patches but differ noticeably across dimensions within each patch. The latent variables capture this pattern well: both $\bm{z\_mean}_{1:N}$ and $\bm{z\_logvar}_{1:N}$ remain stable over patches while preserving significant variation across dimensions. This indicates that the latent representation encodes consistent temporal patterns while maintaining important feature-level distinctions.
Similarly, the generative model’s latent variables learned via variational inference exhibit near-uniformity across patches with meaningful dimension-wise variation. Such a well-structured latent representation, combined with the hidden states $\bm{h}_{1:N}$, offers a comprehensive and informative summary of the time series, thereby supporting accurate and robust forecasting.
As demonstrated in Fig. 13, the variational inference training converges effectively, and the latent states successfully capture the underlying temporal features of the data. This visualization validates both the model’s representational power and the interpretability of its latent space in modeling complex temporal dynamics.
4.2.4 Model Efficiency
We compared the efficiency of different models on two datasets with varying dimensionality: the high-dimensional Solar dataset (137 dimensions) and the lower-dimensional Weather dataset (21 dimensions). The prediction horizon was set to 192 for both datasets. Efficiency metrics include MSE, the number of model parameters (Params), and the theoretical floating-point operations (FLOPs) required for a single forward pass. The results are presented in Fig. 14.
(a) Weather Dataset
(b) Solar Dataset
Figure 14: Comparison of model efficiency on Weather and Solar datasets. The prediction horizon for all models in both datasets is set to 192. Params denotes the number of model parameters, and FLOPs represents the theoretical floating-point operations required for a single forward pass.
The results indicate that xLSTM-based models generally achieve lower MSE values, while their Params and FLOPs lie between those of Linear-based models and Transformer-based models. This demonstrates that xLSTM-based models offer a favorable balance between predictive accuracy and computational efficiency.
For StoxLSTM, the training phase requires an inference model to approximate the posterior distribution of the time series, resulting in substantially higher parameter counts and FLOPs during training. However, during testing, only the generative model is utilized, leading to a significant reduction in both model parameters and computational costs. Notably, StoxLSTM’s generative model achieves high efficiency, as its recurrent units enhance the original xLSTM structure without introducing additional complex modules. Future work will focus on simplifying the inference model to further enhance overall efficiency.
@@ 5 Conclusion
In this paper, we propose StoxLSTM, a stochastic xLSTM architecture. By integrating a designed SSM, StoxLSTM effectively captures the dynamic evolution of time series data. Our model demonstrates a strong capability in both long- and short-term forecasting tasks. Extensive experiments on multiple publicly available benchmark datasets show that StoxLSTM consistently outperforms existing state-of-the-art methods in terms of prediction accuracy and robustness. These results validate the effectiveness of combining the xLSTM framework with the SSM design to enhance temporal modeling.
However, we impose a Gaussian assumption on the latent variables, which may restrict the model’s ability to fully capture highly complex or multimodal temporal dynamics present in real-world time series. Second, due to the VAE-like deep state space model structure, StoxLSTM tends to produce relatively smooth predictions, limiting its capability to accurately model and forecast fine-grained short-term fluctuations and detailed temporal variations. Finally, the training process of StoxLSTM relies on an inference model to approximate the posterior distribution of latent variables, which significantly increases the number of parameters and computational overhead. Future work will focus on relaxing the Gaussian assumption, enhancing the model’s sensitivity to short-term dynamics, and developing more efficient inference mechanisms to reduce computational costs.
6 Acknowledgment
This work was supported in part by the National Natural Science Foundation of China under Grant 62301031, in part by the National Key Laboratory of Science and Technology on Space-Born Intelligent Information Processing under Grant TJ-01-25-01, and in part by the Young Elite Scientists Sponsorship Program by CAST under Grant YESS20240231.